# Near-Field Localization and Phase Shift Optimization for RIS-Assisted Non-Ideal OFDM Systems

### v2

@@ IV Hybrid Cramér-Rao Lower Bound Analysis
In this section, we derive HCRLB to evaluate the joint estimation, in the presence of a random parameter $\bm{\theta}$ in the set of parameters to be estimated. The HCRLB for $\bm{\lambda}$, the parameters to be estimated, is given by [7]
@@
where $\mathbf{B}$ is the Bayesian information matrix (BIM), which can be calculated as
$\mathbf{B}=\mathbb{E}_{\bm{\theta}}\left[\bm{\mathrm{FIM}}(\bm{y};\bm{\lambda})\right]+\mathbb{E}_{\bm{\theta}}\left[-\frac{\partial}{\partial\bm{\lambda}}\left(\frac{\partial\log p(\bm{\theta})}{\partial\bm{\lambda}}\right)^{\top}\right].$ (28)
In order to evaluate the positioning accuracy and the CFO-PN joint MSE, we define a new parameter vector $\bm{\alpha}=\bm{g}(\bm{\lambda})\triangleq\left[\bm{\underline{\gamma}},x_{u},y_{u},z_{u}\right]^{\top}$. In the following subsections, we first derive the expression for matrix $\mathbf{B}$. Then, we derive the transition matrix to transform $\mathbf{B}$ into the HCRLB relative to the parameter vector $\bm{\alpha}$.
@@ IV-A Derivation of $\mathbb{E}_{\bm{\theta}}\left[\bm{\mathrm{FIM}}(\bm{y};\bm{\lambda})\right]$
To derive $\mathbb{E}_{\bm{\theta}}\left[\bm{\mathrm{FIM}}(\bm{y};\bm{\lambda})\right]$, we begin by calculating the Fisher information matrix (FIM) for $\bm{\lambda}$, as expressed by
@@
To calculate (29), the following gives the derivative of $\bm{\mu}$ with respect to each element in vector $\bm{\lambda}$.
$\frac{\partial\bm{\mu}}{\partial\phi}$ $=\sqrt{P}\bm{\Lambda}\bm{\Lambda}_{\phi}\bm{\Lambda}_{\bm{\theta}}\mathbf{S}^{\top}\bm{h},$ (30a)
$\frac{\partial\bm{\mu}}{\partial\bm{\theta}[n]}$ $=\sqrt{P}{\rm Diag}(\bm{a}[n])\bm{\Lambda}_{\phi}\mathbf{S}^{\top}\bm{h},$ (30b)
$\frac{\partial\bm{\mu}}{\partial d_{Ru}}$ $=\sqrt{P}\bm{\Lambda}_{\phi}\bm{\Lambda}_{\bm{\theta}}\mathbf{S}^{\top}\frac{\partial\bm{h}}{\partial d_{Ru}},$ (30c)
$\frac{\partial\bm{\mu}}{\partial\phi_{az}}$ $=\sqrt{P}\bm{\Lambda}_{\phi}\bm{\Lambda}_{\bm{\theta}}\mathbf{S}^{\top}\frac{\partial\bm{h}}{\partial\phi_{az}},$ (30d)
@@
where $\bm{\Lambda}$ is a diagnal matrix, $\left[\bm{\Lambda}\right]_{n,n}=\frac{j2\pi n}{N}$, $\bm{a}[n]=[\bm{0}_{1\times n},je^{j\bm{\theta}[n]},\bm{0}_{1\times(N-n-1)}]^{\top}$.
It is not difficult to observe that in (30), the terms related to $\phi$ and $\bm{\theta}$ all appear in the exponential terms of the diagonal elements of the corresponding matrix. Hence, they would disappear when calculating $\frac{\partial\bm{\mu}{\rm{}^{\rm H}}}{\partial\lambda_{i}}\frac{\partial\bm{\mu}}{\partial\lambda_{j}}$ in (29). That is to say, the calculated FIM is not related to parameter $\bm{\theta}$, and
@@
Since $\bm{\theta}\sim\mathcal{N}(\bm{0}_{N\times 1},\bm{\Psi})$, we can calculate $\log p(\bm{\theta})$ by $\log p(\bm{\theta})=-\frac{1}{2}\bm{\theta}^{\top}\bm{\Psi}^{-1}\bm{\theta}-\frac{N}{2}\log(2\pi)-\frac{1}{2}\log\det\bm{\Psi}$, and its partial to $\bm{\lambda}$ can be calculated by
@@ IV-C Derivation of the Transformed HCRLB
Due to the ambiguity between the parameters to be estimated, the MSE of CFO and PN is jointly calculated using (26), and when computing the position error bound (PEB), the position parameters are required to be in the form of 3D Cartesian coordinates $x_{u}$, $y_{u}$, and $z_{u}$. Therefore, the parameter vector $\bm{\lambda}$ needs to be transformed into $\bm{\alpha}$. The HCRLB for $\bm{\alpha}$ is given by [9]
@@
where for any UE position $\bm{p}_{u}$, $\mathbf{\Xi}(\bm{p}_{u})$ is the transition matrix defined as
@@
After obtaining $\mathbf{HCRLB}(\bm{p}_{u},\bm{w})$, $\rm PEB$ for position estimation and $\rm HCRLB_{CFO-PN}$ for CFO-PN estimation can be calculated as
${\rm PEB}=\sqrt{{\rm tr}([\mathbf{HCRLB}(\bm{p}_{u},\bm{w})]_{N:N+2,N:N+2})},$ (38)
@@
Both ${\rm PEB}$ and ${\rm HCRLB_{CFO-PN}}$ are related to UE’s location $\bm{p}_{u}$ and RIS phase shift $\bm{w}$.
V RIS Phase Shift Optimization
In this section, our goal is to optimize the phase shift of the RIS to enhance theoretical positioning accuracy, i.e., PEB. According to equation (38), minimizing the PEB at a specific location requires knowledge of that location. However, this is not possible in practical positioning scenarios. To address this issue, we select multiple sampling points within the AOI randomly, then calculate and minimize their average PEB. Accordingly, the optimization problem for RIS phase shift can be expressed as
@@ V RIS Phase Shift Optimization
Problem (V) is non-convex because the objective function does not exhibit convexity with respect to vector $\bm{w}$. We define $\mathbf{W}\triangleq\bm{\bar{w}}\bm{\bar{w}}^{\rm H}$, where $\bm{\bar{w}}\triangleq\bm{w}^{*}$, which needs to satisfy $\mathbf{W}\succcurlyeq 0$ and ${\rm rank}(\mathbf{W})=1$. Since the rank-1 constraint is non-convex, we employ the SDR method to relax this constraint. Consequently, the problem (V) can be transformed into
@@
Inspired by [12], we utilize the property of the Schur complement [6], and the problem can be expressed in the form of SDP. For each calculated PEB at a particular position $\bm{p}_{u}$, we define an auxiliary matrix $\mathbf{Z}_{u}$ that satisfies the following expression
@@
where $\tilde{\bm{\Xi}}(\bm{p}_{u})\triangleq\mathbf{T}\mathbf{\Xi}(\bm{p}_{u})$, and $\mathbf{T}\triangleq[\bm{0}_{N\times N}\quad\mathbf{I}_{3}]\in\mathbb{R}^{(N+3)\times(N+3)}$ can extract the PEB portion from the HCRLB.
The optimization problem in (V) can be converted to
@@
$\begin{array}[]{ll}s.t.&\left[\mathbf{W}\right]_{r,r}=1,r=1,\cdots,N_{R},\\ &\mathbf{W}\succcurlyeq 0,\\ &\displaystyle\frac{1}{U}\sum_{u=1}^{U}\begin{bmatrix}\mathbf{Z}_{u}&\tilde{\bm{\Xi}}(\bm{p}_{u})\\ \tilde{\bm{\Xi}}^{\top}(\bm{p}_{u})&\mathbf{B}(\bm{p}_{u},\mathbf{W})\end{bmatrix}\succcurlyeq 0.\end{array}$ (48)
Problem (V) is a convex SDP problem, and can be optimally solved using established convex optimization solvers, such as CVX [30]. Since there is no rank-1 constraint in the optimization problem, it is highly likely that the obtained matrix $\mathbf{W}$ will have a rank greater than one. Therefore, decomposing it into the desired RIS phase shift vector $\bm{w}$ is a crucial subsequent problem. As in [29], we first perform an eigenvalue decomposition on the optimal $\mathbf{W}$, then take the largest eigenvector and multiply it by the square root of the largest eigenvalue and normalize the vector, obtaining $\bm{\bar{w}}$. Finally, we calculate the RIS phase shift by $\bm{w}=\bm{\bar{w}}^{*}$. The decomposition process can also be accomplished using the Gaussian randomization method as in [28].
In order to meet the requirements of (V), we next derive the expression of $\mathbf{B}$ that is related to $\mathbf{W}$. We first rewrite (4) as
@@
which separates $\bm{w}$ from $\bm{h}$, with $\mathbf{G}\in\mathbb{C}^{N_{R}\times N}$ and $\left[\mathbf{G}\right]_{r,n}=\sqrt{\rho_{aR,n,r}}\sqrt{\rho_{Ru,n,r}}e^{-j2\pi\frac{nB}{N}(\tau_{ar}+\tau_{ru})}$.
@@
According to (28), only the $\bm{\mathrm{FIM}}(\bm{y};\bm{\lambda})$ part is related to $\bm{w}$, hence we only need to derive the relationship of this part of $\mathbf{B}$ and $\mathbf{W}$. We can rewrite the FIM in a summation form:
The partial derivatives of $\bm{\mu}[n]$ and $\bm{\mu}[n]^{*}$ with respect to all parameters to be estimated are derived in Appendix A. It is observed that the expressions for $\frac{\partial\bm{\mu}[n]^{*}}{\partial\lambda_{i}}$ all end with $\bm{\bar{w}}$, while the expressions for $\frac{\partial\bm{\mu}[n]}{\partial\lambda_{j}}$ all start with $\bm{\bar{w}}^{\rm H}$ for the vector parts. Therefore, when calculating the FIM, we can easily obtain the expression with respect to $\mathbf{W}$.
VI Numerical Results

### v3

@@ IV Hybrid Cramér-Rao Lower Bound Analysis
The HCRLB for the joint estimation of the CFO, PN, and UE position is derived in this section. This helps quantify the performance lower bound of the proposed AO-based estimation algorithm, and also serves as the objective function for optimizing the RIS phase shifts, $\bm{w}$, in the next section. The HCRLB for the estimation of $\bm{\lambda}$ is given by [34, Eq. (4.610)]
@@
where $\mathbf{B}$ is the Bayesian information matrix (BIM), which can be given by [34, Eq. (4.611)]
$\mathbf{B}=\mathbb{E}_{\bm{\theta}}\left[\bm{\mathrm{FIM}}(\bm{y};\bm{\lambda})\right]+\mathbb{E}_{\bm{\theta}}\left[-\frac{\partial}{\partial\bm{\lambda}}\left(\frac{\partial\log p(\bm{\theta})}{\partial\bm{\lambda}}\right)^{\top}\right],$ (28)
where $\bm{\mathrm{FIM}}(\bm{y};\bm{\lambda})$ stands for the Fisher’s information matrix (FIM) for $\bm{\lambda}$ under the observation of received signal $\bm{y}$ at the UE, and $\mathbb{E}_{\bm{\theta}}\left[-\frac{\partial}{\partial\bm{\lambda}}\left(\frac{\partial\log p(\bm{\theta})}{\partial\bm{\lambda}}\right)^{\top}\right]$ is the prior information matrix. Interested readers may refer to [34] for more details.
@@ IV-A Derivation of $\mathbb{E}_{\bm{\theta}}\left[\bm{\mathrm{FIM}}(\bm{y};\bm{\lambda})\right]$
To derive $\mathbb{E}_{\bm{\theta}}\left[\bm{\mathrm{FIM}}(\bm{y};\bm{\lambda})\right]$ in (28), we first analyze the FIM for $\bm{\lambda}$, as given by
@@
To calculate (29), the first-order derivative of $\bm{\mu}$ with respect to each element in vector $\bm{\lambda}$ is derived as follows.
$\frac{\partial\bm{\mu}}{\partial\phi}$ $=\sqrt{P}\bm{\Lambda}\bm{\Lambda}_{\phi}\bm{\Lambda}_{\bm{\theta}}\mathbf{S}^{\top}\bm{h};$ (30a)
$\frac{\partial\bm{\mu}}{\partial\bm{\theta}[n]}$ $=\sqrt{P}{\rm Diag}(\bm{a}[n])\bm{\Lambda}_{\phi}\mathbf{S}^{\top}\bm{h};$ (30b)
$\frac{\partial\bm{\mu}}{\partial d_{Ru}}$ $=\sqrt{P}\bm{\Lambda}_{\phi}\bm{\Lambda}_{\bm{\theta}}\mathbf{S}^{\top}\frac{\partial\bm{h}}{\partial d_{Ru}};$ (30c)
$\frac{\partial\bm{\mu}}{\partial\phi_{az}}$ $=\sqrt{P}\bm{\Lambda}_{\phi}\bm{\Lambda}_{\bm{\theta}}\mathbf{S}^{\top}\frac{\partial\bm{h}}{\partial\phi_{az}};$ (30d)
@@
where $\bm{\Lambda}$ is a diagnal matrix, $\left[\bm{\Lambda}\right]_{n,n}=\frac{j2\pi n}{N}$, and $\bm{a}[n]=[\bm{0}_{1\times n},je^{j\bm{\theta}[n]},\bm{0}_{1\times(N-n-1)}]^{\top}$.
It is not difficult to see that in (30), only $\bm{\Lambda}_{\phi}$, $\bm{\Lambda}_{\bm{\theta}}$ and ${\rm Diag}(\bm{a}[n])$ are related to $\phi$ and $\bm{\theta}$, which are all diagonal matrices. Additionally, both $\phi$ and $\bm{\theta}$ appear in the exponent of diagonal elements in these matrices. Hence, they can be suppressed when calculating $\frac{\partial\bm{\mu}{\rm{}^{\rm H}}}{\partial\lambda_{i}}\frac{\partial\bm{\mu}}{\partial\lambda_{j}}$ in (29). In other words, the calculated FIM does not rely on the parameter $\bm{\theta}$, i.e.,
@@
Take $\frac{\partial\bm{\mu}^{\rm H}}{\partial d_{Ru}}\frac{\partial\bm{\mu}}{\partial d_{Ru}}$ as an example. We have
$\frac{\partial\bm{\mu}^{\rm H}}{\partial d_{Ru}}\frac{\partial\bm{\mu}}{\partial d_{Ru}}$ $=\sqrt{P}\frac{\partial\bm{h}^{\rm H}}{\partial d_{Ru}}\mathbf{S}^{*}\bm{\Lambda}_{\bm{\theta}}^{\rm{H}}\bm{\Lambda}_{\phi}^{\rm H}\sqrt{P}\bm{\Lambda}_{\phi}\bm{\Lambda}_{\bm{\theta}}\mathbf{S}^{\top}\frac{\partial\bm{h}}{\partial d_{Ru}}$ (32)
$=P\frac{\partial\bm{h}^{\rm H}}{\partial d_{Ru}}\mathbf{S}^{*}\mathbf{S}^{\top}\frac{\partial\bm{h}}{\partial d_{Ru}}$
which is unrelated to $\phi$ or $\bm{\theta}$, since $\bm{\Lambda}_{\phi}^{\rm H}\bm{\Lambda}_{\phi}=\bm{\Lambda}_{\bm{\theta}}^{\rm{H}}\bm{\Lambda}_{\bm{\theta}}=\mathbf{I}_{N}$. This can be extended to $\frac{\partial\bm{\mu}{\rm{}^{\rm H}}}{\partial\lambda_{i}}\frac{\partial\bm{\mu}}{\partial\lambda_{j}},\forall i,j$, and is omitted here.
@@
Since $\bm{\theta}\sim\mathcal{N}(\bm{0}_{N\times 1},\bm{\Psi})$, we can calculate $\log p(\bm{\theta})$ as $\log p(\bm{\theta})=-\frac{1}{2}\bm{\theta}^{\top}\bm{\Psi}^{-1}\bm{\theta}-\frac{N}{2}\log(2\pi)-\frac{1}{2}\log\det\bm{\Psi}$, and its partial derivative with respect to $\bm{\lambda}$ can be calculated by
@@ IV-C Derivation of the Transformed HCRLB
So far, we have derived the expression for $\mathbf{B}$, which is an HCRLB matrix that gives analytical estimation accuracy lower bounds for the parameters in $\bm{\lambda}$. However, when assessing the proposed estimation algorithm, we adopt the RMSE and CFO-PN joint MSE metrics introduced in Section III-E. Hence, we define a new parameter vector $\bm{\alpha}=\bm{g}(\bm{\lambda})\triangleq\left[\bm{\underline{\gamma}},x_{u},y_{u},z_{u}\right]^{\top}$, corresponding to the two metrics. Then, $\mathbf{B}$ can be transformed into the $\bm{\alpha}$-related HCRLB matrix by [35]
@@
where, for any UE position $\bm{p}_{u}$, $\mathbf{\Xi}(\bm{p}_{u})$ is the transition matrix defined as
@@
After obtaining $\mathbf{HCRLB}(\bm{p}_{u},\bm{w})$, the position error bound (${\rm PEB}$) for evaluating the UE position estimation and $\rm HCRLB_{CFO-PN}$ for evaluating the CFO-PN joint estimation can be given by
${\rm PEB}=\sqrt{{\rm tr}([\mathbf{HCRLB}(\bm{p}_{u},\bm{w})]_{N:N+2,N:N+2})};$ (39)
@@
Both ${\rm PEB}$ and ${\rm HCRLB_{CFO-PN}}$ depend on the UE’s position, $\bm{p}_{u}$, and the RIS phase shifts, $\bm{w}$.
V RIS Configuration
In this section, we further configure the phase shifts $\bm{w}$ of the RIS to minimize the analytical lower bound of the localization accuracy, i.e., PEB. According to (39), the PEB depends on the UE position, which indicates that minimizing the PEB at a specific UE position requires knowledge of that position. However, this is not possible in practice. To address this issue and enhance the algorithm’s robustness, we take the average PEB within the area of interest (AOI) as the objective function. Specifically, we randomly sample $U$ positions within the AOI and calculate their average PEB. Accordingly, the problem of interest is formulated as
@@ V RIS Configuration
Problem (V) is not convex because of the non-convexity of the objective function with respect to vector $\bm{w}$, as well as the unit-module constraint.
To convexify (V), we first define $\mathbf{W}\triangleq\bm{\bar{w}}\bm{\bar{w}}^{\rm H}$, where $\bm{\bar{w}}\triangleq\bm{w}^{*}$, which satisfies $\mathbf{W}\succcurlyeq 0$ and ${\rm rank}(\mathbf{W})=1$. Since the rank-1 constraint is non-convex, we employ the SDR method to relax this constraint. Consequently, the problem in (V) can be transformed into
@@
According to (35) and (39), the optimization problem in (V) remains non-convex because of the complex objective function. To transform (V) into the convex SDP form, for each calculated PEB at a particular position $\bm{p}_{u}$, we define an auxiliary matrix $\mathbf{Z}_{u}$ that satisfies
@@
where $\tilde{\bm{\Xi}}(\bm{p}_{u})\triangleq\mathbf{T}\mathbf{\Xi}(\bm{p}_{u})$, and $\mathbf{T}\triangleq[\bm{0}_{N\times N}\quad\mathbf{I}_{3}]\in\mathbb{R}^{(N+3)\times(N+3)}$ can extract the PEB from the HCRLB. According to the property of the Schur complement and the positive semidefiniteness of $\mathbf{B}(\bm{p}_{u},\mathbf{W})$, (45) can be transformed to
$\begin{bmatrix}\mathbf{Z}_{u}&\tilde{\bm{\Xi}}(\bm{p}_{u})\\ \tilde{\bm{\Xi}}^{\top}(\bm{p}_{u})&\mathbf{B}(\bm{p}_{u},\mathbf{W})\end{bmatrix}\succcurlyeq 0.$ (46)
Thus, the problem in (V) can be converted to
@@
$\begin{array}[]{ll}s.t.&\left[\mathbf{W}\right]_{r,r}=1,r=1,\cdots,N_{R},\\ &\mathbf{W}\succcurlyeq 0,\\ &\displaystyle\frac{1}{U}\sum_{u=1}^{U}\begin{bmatrix}\mathbf{Z}_{u}&\tilde{\bm{\Xi}}(\bm{p}_{u})\\ \tilde{\bm{\Xi}}^{\top}(\bm{p}_{u})&\mathbf{B}(\bm{p}_{u},\mathbf{W})\end{bmatrix}\succcurlyeq 0,\end{array}$ (50)
where, according to the relationship in (45) and the calculation of PEB in (39), minimizing the objective function results in minimizing the average square of PEB, which also minimizes the initial objective function, the average PEB.
Proposition 1. $\mathbf{B}(\bm{p}_{u},\mathbf{W})$ is linear to $\mathbf{W}$.
Proof. According to (28), only $\bm{\mathrm{FIM}}(\bm{y};\bm{\lambda})$ depends on $\bm{w}$ in $\mathbf{B}$. Hence, we only need to prove the linearity of $\bm{\mathrm{FIM}}(\bm{y};\bm{\lambda})$ to $\mathbf{W}$, which can be rewritten as
To derive the explicit expressions of $\frac{\partial\bm{\mu}[n]^{*}}{\partial\lambda_{i}},\forall i$ and $\frac{\partial\bm{\mu}[n]}{\partial\lambda_{j}},\forall j$ with respect to $\bar{\bm{w}}$, we rewrite (6) as
@@
with $\mathbf{G}\in\mathbb{C}^{N_{R}\times N}$ and $\left[\mathbf{G}\right]_{r,n}=\sqrt{\rho_{aR,n,r}}\sqrt{\rho_{Ru,n,r}}e^{-j2\pi\frac{nB}{N}(\tau_{ar}+\tau_{ru})}$.
@@
The partial derivatives of $\bm{\mu}[n]$ and $\bm{\mu}[n]^{*}$ with regard to all parameters to be estimated are derived in Appendix A. It can be observed that $\bm{\bar{w}}$ is only in the last term of $\frac{\partial\bm{\mu}[n]^{*}}{\partial\lambda_{i}},\forall i$, while $\bm{\bar{w}}^{\rm H}$ appears as the first term of $\frac{\partial\bm{\mu}[n]}{\partial\lambda_{j}},\forall j$, for the vector parts. Hence, the product of $\frac{\partial\bm{\mu}[n]^{*}}{\partial\lambda_{i}},\forall i$ and $\frac{\partial\bm{\mu}[n]}{\partial\lambda_{j}},\forall j$ would lead to $\mathbf{W}$. As a result, as can be seen from (51), every term of $\bm{\mathrm{FIM}}(\bm{y};\bm{\lambda})$ is linear to $\mathbf{W}$ because the operations of taking the real parts, summation, and scalar multiplication preserve linearity. Hence, $\mathbf{B}(\bm{p}_{u},\mathbf{W})$ is linear to $\mathbf{W}$. ∎
According to Proposition 1, (V) is a convex SDP problem, and can be optimally solved using established convex optimization solvers, such as off-the-shelf CVX toolboxes [36]. Since the rank-1 constraint is relaxed via the SDR process, the obtained matrix $\mathbf{W}$ is likely to have a rank greater than 1. Therefore, decomposing it into the desired RIS phase shift vector $\bm{w}$ is a crucial subsequent problem. As done in [37], we first perform EVD on the optimal $\mathbf{W}$, then take the largest eigenvector and multiply it with the square root of the largest eigenvalue and normalize the vector, obtaining $\bm{\bar{w}}$. Finally, we calculate the RIS phase shifts by $\bm{w}=\bm{\bar{w}}^{*}$. The decomposition process can also be accomplished using the Gaussian randomization method, as done in [38].
Being an SDP, the problem in (V) can be solved using the interior point method in the CVX toolboxes. Thus, the convergence of the proposed RIS configuration is guaranteed. The computational complexity of the proposed RIS configuration is primarily influenced by the interior point method employed to solve (V). The worst-case complexity of the interior point method is $\mathcal{O}(\max\left\{N_{con},N_{var}\right\}^{4}\sqrt{N_{v}})$, where $N_{con}$ and $N_{var}$ are the constraint and variable numbers, respectively [39]. For the problem in (V), ${N_{con}}=N_{R}+2$ and ${N_{var}}=N_{R}^{2}+U(N+3)^{2}$. Suppose that the convergence accuracy is $\epsilon$. The complexity of solving (V) is $\mathcal{O}((N_{R}^{2}+UN^{2})^{4.5}\log\frac{1}{\epsilon})$.
VI Numerical Evaluation